1) Samson Company had the following balances and transactions during 2013.

Beginning inventory

10 units at $70

June 10

Purchased 20 units at $80

The company sold 23 units in 2013

What would the company's Cost of goods sold be on the December 31, 2013 income statement if the Last-In, First-Out costing method is used? (Answers are rounded to the nearest dollar.)

A) $1,740

B) $1,810

C) $1,690

D) $1,540

a) What would the company's Inventory amount be on the December 31, 2013 balance sheet if the Last-In, First-Out costing method is used?

A) $490

B) $554

C) $560

D) $540

b) What would the company's Cost of goods sold be on the December 31, 2013 income statement if the First-In, First-Out costing method is used? (Answers are rounded to the nearest dollar.)

A) $1,740

B) $1,810

C) $1,690

D) $1,540

c) What would the company's Inventory amount be on the December 31, 2013 balance sheet if the First-In, First-Out costing method is used?

A) $490

B) $554

C) $560

D) $540
Answer #1

What would the company's Cost of goods sold be on the December 31, 2013 income statement if the Last-In, First-Out costing method is used?

      b.) $1810 ( in fifo method, last purchased goods is sold first. So, cost of goods sold = 20*80+3*70 =1810)

a) What would the company's Inventory amount be on the December 31, 2013 balance sheet if the Last-In, First-Out costing method is used?

     a.) $490. ( As per LIFO, stock from opening inventory is left for 7 units @ 70.So, value of closing inventory=7*70)

b) What would the company's Cost of goods sold be on the December 31, 2013 income statement if the First-In, First-Out costing method is used?

     a.) $1740. (As per FIFO, first purchased goods is sold first. Cost of goods sold =10*70 + 13*80 =1740)

c) What would the company's Inventory amount be on the December 31, 2013 balance sheet if the First-In, First-Out costing method is used?

    c.) $560 (As per FIFO, first purchased goods is sold first. Cost of inventory left =7*80 =560)